# Eigenvalue Gaps of the Laplacian of Random Graphs

### v1

@@ 3. Supporting lemmas
We let $I:=[-ck/\sqrt{n},ck/\sqrt{n}]$. Let us assume that $\lambda_{i}\in I$ and ${\mathbf{v}}_{i}$ is such that $(L_{n}+F){\mathbf{v}}_{i}=\lambda_{i}{\mathbf{v}}_{i}$. As $L_{n}+F$ is symmetric, and so the eigenvectors $v_{i}$ are orthogonal, then
$\left\|(L_{n}+F){\mathbf{v}}_{i}\right\|_{2}\leq\frac{ck}{\sqrt{n}}$ (11)
Write ${\mathbf{v}}_{j}=(v_{j1},\ldots,v_{jn})^{T}$ and let ${\mathbf{c}}_{i}$ denote the $i^{\text{th}}$ column of $L_{n}+F$. Let $V$ be the $k\times n$ matrix formed by the row vectors ${\mathbf{v}}_{i}^{T}$ and let ${\mathbf{w}}_{1},\dots,{\mathbf{w}}_{n}$ be its columns. Clearly, (11) is equivalent to
@@
For $J\subset[n]$, we use $V_{J}$ to indicate the matrix formed by the columns ${\mathbf{w}}_{j}$ with $j\in J$. Oberve that by construction,
@@
We now show that there exists a well-conditioned minor of submatrix of $V$. By Lemma 3.1, for any $11\leq l\leq k-$ 1, there exist distinct indices $i_{1},\dots,i_{l}$ such that
@@
Let $H$ denote the subspace spanned by $({\mathbf{c}}_{i_{l+1}},\ldots,{\mathbf{c}}_{i_{n}})$. (14) implies that
@@
Thus, we have reduced the event in (11) to the even that
${\operatorname{dist}}({\mathbf{c}}_{i_{1}},H)=O\left(\frac{k^{2}}{(k-l)}\right)\wedge\cdots\wedge{\operatorname{dist}}({\mathbf{c}}_{i_{l}},H)=O\left(\frac{k^{2}}{(k-l)}\right)$
@@
Note that now we have that ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is independent of ${\mathbf{c}}_{j,\{l,\dots,n\}}$ for $i,j\leq l$ and $i\neq j$. However, due to the structure of the Laplacian, ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is not independent of $H_{l,\dots,n}$
We now set $l=k/2$ and we introduce an integer parameter $\tau$, which will be a sufficiently large constant depending on $c$ and $p$, and divide our indices $1,\dots,l$ into $l/2\tau$ sets,
$J_{1}=\{1,\dots,2\tau\},J_{2}=\{2\tau+1,4\tau\},\dots,J_{l/2\tau}=\{l-2\tau+1,\dots,l\}$
@@
Every coordinate of $\bar{{\mathbf{d}}}_{i,\{l,\dots,n\}}:={\mathbf{d}}_{i,\{l,\dots,n\}}^{*}+{\mathbf{d}}_{i,\{l,\dots,n\}}^{**}$ is an i.i.d. binomial random variable, For a random variable $X\sim B(2\tau,p)$, by Chernoff’s bound, for any $10<\delta<$ 1,
@@
Now for $i\in[l/2\tau]$, we introduce the following auxiliary randomness inspired again by the “switching” method. We define a new random variable $\hat{{\mathbf{c}}}^{*}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{*})_{j}=({\mathbf{c}}^{*}_{i})_{j}$ for $j\notin D_{i}$ and for $j\in D_{i}$,
$(\hat{{\mathbf{c}}}^{*}_{i})_{j}=\begin{cases}&(\hat{{\mathbf{c}}}_{i}^{*})_{j}\text{ with probability }1/2\\ &({\mathbf{d}}_{i}^{**})_{j}+{\mathbf{f}}_{i}^{*}\text{ with probability }1/2\end{cases}$
and similarly, we define a new random variable $\hat{{\mathbf{c}}}^{**}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{**})_{j}=({\mathbf{c}}^{**}_{i})_{j}$ for $j\notin D_{i}$ and for $j\in D_{i}$,
@@
We observe that $\hat{{\mathbf{c}}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$ has the same joint distribution as ${\mathbf{c}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$. Furthermore, the key benefit of this construction is that even upon conditioning on the outcome of $H_{l,\dots,n}$, ${\mathbf{c}}_{i}$, and ${\mathbf{c}}_{i}^{*}$, $\hat{{\mathbf{c}}}_{i}^{*}$ is sufficiently random with independent entries for $j\in I_{i}$.
@@
On the event $\mathcal{E}$, we have at least $l/2\tau\alpha$ good indices. For any good index $i$, by (15),
${\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{l,\dots,n})\leq O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)$ $\leq{\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{D_{i}})=O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)$
$\leq C\exp(-\Theta(k))$
where the last inequality follows from Lemma 4.1 and $k$ sufficiently large. Note that our choice of $l$ and size of $|D_{i}|$ guarantees that the corank of $H_{D_{i}}$ is at least $k/3$. Therefore, by independence of the randomness on the good indices, we have that
@@
Proof.
(of Lemma 4.1) We show that $f({\mathbf{x}})=\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is convex and 1-Lipschitz. For convexity, we observe that for $0\leq s\leq 1$,
@@
Next, note that ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{n}$,
@@
${\mathbb{P}}(|f({\mathbf{x}})-M(f({\mathbf{x}})|\geq\lambda T)\leq K\exp(-\kappa\lambda^{2})$ (18)
@@
$=\left(T^{4}k+T^{2}\|{\mathbf{u}}\|_{2}^{2}k\right)$
@@
Case 1. We assume that $(1-2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}\leq\|{\mathbf{u}}\|_{2}\leq(1+2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}$, then we have that
@@
By Markov’s inequality, the median of $|Y|$ is $O(T^{2}\sqrt{k})$, which implies that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}$ is at least $\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}+\|u\|_{2}^{2}-O(T^{2}\sqrt{k})\geq T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})$. We can therefore deduce that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is at least $\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})}=\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}}-O(T^{2})$. Returning to Talagrand’s inequality, (18), we find that
@@
Case 2. Now we consider the case where $\|{\mathbf{u}}\|_{2}\leq(1-2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|{\mathbf{u}}\|_{2}\geq(1+2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$. Then, if $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}\leq t\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ then by the triangle inequality, either $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\leq(1-t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\geq(1+t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ which reduces to Case 1. ∎
5. No-gaps delocalization: proof of Theorem 2.3 for the approximated vectors
First, recall the Erdős-Littlewood-Offord result from [NW24].
@@
  • Assume that at least $N^{\prime}$ elements of the $w_{i}$ satisfy $|w_{i}|\geq r$. Let $X=(x_{1},\dots,x_{N})$ where $x_{i}$ are iid copies of a random variable $\xi$ of mean zero, variance one, and bounded $(2+\varepsilon)$-moment. Then
@@
  • Furthermore, if we assume that for all $w\in{\mathbb{R}}$, the vector $\operatorname{supp}({\mathbf{w}}-w\mathbf{1})$ has at least $N^{\prime}$ components that have absolute values at least $r$. Then for $X=(x_{1},\dots,x_{N})$ with $x_{i},$ for $1\leq i\leq N$, being i.i.d. copies of $\xi$ we have the following affine analog
@@
$\frac{1}{\log\log n}\leq\lambda_{n}<1/2.$
It what follows we will mainly work with $\overline{L}_{n}$, the proof for $\overline{L}^{\prime}_{n}$ (or for $\overline{L}^{\prime}_{n-1}$) is identical.
@@ 5.1. Decomposition of vectors
First, we let $I$ be the set of first $\lambda_{n}n$ indices (in the end we will take union bound over $I$) and will bound the probability of the following event $\mathcal{E}$: there exists a ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) whose first $\lceil\lambda_{n}n\rceil$ coordinates are the same (in the $l_{2}$-norm), that is for some $a$
@@
We let $t_{0}>1$ be a jumping parameter to be chosen later, for instance we can choose
@@
Given ${\mathbf{v}}$, we will partition $[n]$ into two subsets, $I_{m}$ (mixed) and $I_{s}$ (sparse), where $I_{s}$ is the union of the intervals $J_{i_{0}},J_{{i_{1}}},\dots$ above which can be defined step by step below: at each step $i\geq 0$, find a ${\mathbf{v}}_{J_{i}}$ such that there exists $a_{i}\in{\mathbb{Z}}/D$ and a subset $J_{i}^{\prime}\subset J_{i}$ of size at least $|J_{i}|-\alpha_{n}\lambda_{n}n$ which satisfies
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}\leq\delta_{n}t_{0}^{i}.$ (25)
Hence $I_{s}$ is associated with a maximum index $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ and a partial ordering of $\{1,\dots,\lfloor\lambda_{n}^{-1}\rfloor\}$ where the $J_{i}$ appeared in each step $i$.
@@
Each ${\mathbf{v}}$ will also gives rise to a parameter $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ such that all of the intervals from $I_{s}$ are $\delta_{n}t_{0}^{{i_{0}}}$-closed to some $a_{i}\mathbf{1}/\sqrt{n}$, while in $I_{m}$, if $J_{i}$ does not belong to $I_{s}$ then by definition, for any $a\in{\mathbb{Z}}/D$ and any $J_{i}^{\prime}\subset J_{i}$ such that $|J_{i}^{\prime}|\geq|J_{i}|-\alpha_{n}\lambda_{n}n$ we have
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}>\delta_{n}t_{0}^{i_{0}+1}.$
@@
In particularly, Condition 21 of Theorem 5.1 is satisfied, and so
$\rho_{L,\frac{\delta_{n}t_{0}^{i_{0}+1}\lambda_{n}^{-1/2}}{\sqrt{n}}}({\mathbf{v}}_{J_{i}})=O(\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}}).$ (26)
For convenience set
@@
with a saving factor $t_{0}^{1/2}$ for later use.
@@
So for each fixed $a$, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n}\}$ trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size 1.
As a consequence, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n},\mbox{ for some }a\in{\mathbb{Z}}/D,|a|\leq\sqrt{n}\}$, trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size $O(D\sqrt{n})$.
We let $I_{m}=[n]\setminus I_{s}$. We write ${\mathbf{v}}_{m}:={\mathbf{v}}_{I_{m}}$ and ${\mathbf{v}}_{s}:={\mathbf{v}}_{I_{s}}$. Note that ${\mathbf{v}}_{s}$ can be $r_{n}\sqrt{|I_{s}|}$-approximated by a vector ${\mathbf{v}}_{s}^{\prime}$ which is a concatenation of vectors of form $(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}$. We let ${\mathbf{v}}^{\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{m}^{\prime})$ where ${\mathbf{v}}_{m}^{\prime}={\mathbf{v}}_{m}$, and also, basing on Subsection 2.4, without loss of generality we can assume the entries of ${\mathbf{v}}_{m}^{\prime}$ (which is the subvector ${\mathbf{v}}_{J_{i}}$ where there is no $a_{i}$, and from the subvectors ${\mathbf{v}}_{J_{i}\backslash J_{i}^{\prime}}$ where there exists $a_{i}$) are from ${\mathbb{Z}}/D$.
Note that by defition,
@@
where $\sqrt{n}\log n$ comes from the upper bound for $\|L-\lambda_{0}\|_{2}$, and by triangle inequality
@@
As of this point, roughly speaking we have extended the special vector ${\mathbf{v}}_{I}$ to a longest possible vector ${\mathbf{v}}_{s}$ whose entries take only a few values, while ${\mathbf{v}}_{m}$ is the left-over that we no longer detect entries with high multiplicities.
Case 1. Let $\mathcal{E}_{1}$ be the event that there exists a non-zero ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) whose first $\lceil\lambda_{n}n\rceil$ coordinates are the same, and the resulting $I_{m}$ (from ${\mathbf{v}}$) is empty in the sense that for all $i$ we can approximate ${\mathbf{v}}_{J_{i}}$. It is crucial to note that in this case we don’t have to take union bound over $I$ (of Theorem 2.3) because the event $I_{m}$ being empty does not depend on $I$.
@@
We will consider a $\delta_{n}t_{0}^{\lambda_{n}^{-1}}$-approximation of the vectors. There are at most $D^{\alpha_{n}n}\times D^{\lambda_{n}^{-1}}$ choices of ${\mathbf{v}}^{\prime}$ that can arise with empty $I_{m}$ with the first factor stands for the amount of choices of the outliers and the second factors to be the choices of the representation of each interval (assuming that $1/D$ is small compared to $\delta_{n}t_{0}^{\lambda_{n}^{-1}}$, which is necessary for the approximation to make sense).
@@
Then either the smallest or largest values of $a_{i}$ is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the rest. Assume without loss of generality that $a$, the multiplicity of the first segment, is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the multiplicities $j_{0}$, since the level set comes together with a partial order and $j_{0}$ covers every others. Consider the event

### v2

@@ 3. Supporting Lemmas
We let $I:=[-ck/\sqrt{n},ck/\sqrt{n}]$. Let us assume that $\lambda_{i}\in I$ and ${\mathbf{v}}_{i}$ is such that $(L_{n}+F){\mathbf{v}}_{i}=\lambda_{i}{\mathbf{v}}_{i}$. As $L_{n}+F$ is symmetric, the eigenvectors $v_{i}$ are orthogonal, so that
$\left\|(L_{n}+F){\mathbf{v}}_{i}\right\|_{2}\leq\frac{ck}{\sqrt{n}}.$ (11)
Write ${\mathbf{v}}_{j}=(v_{j1},\ldots,v_{jn})^{T}$, and let ${\mathbf{c}}_{i}$ denote the $i^{\text{th}}$ column of $L_{n}+F$. Let $V$ be the $k\times n$ matrix formed by the row vectors ${\mathbf{v}}_{i}^{T}$ and let ${\mathbf{w}}_{1},\dots,{\mathbf{w}}_{n}$ be its columns. Clearly, (11) is equivalent to
@@
For $J\subset[n]$, we use $V_{J}$ to indicate the matrix formed by the columns ${\mathbf{w}}_{j}$ with $j\in J$. Observe that by construction,
@@
We now show that there exists a well-conditioned minor (submatrix) of $V$. By Lemma 3.1, for any $11\leq l\leq k-$ 1, there exist distinct indices $i_{1},\dots,i_{l}$ such that
@@
Let $H$ denote the subspace spanned by $({\mathbf{c}}_{i_{l+1}},\ldots,{\mathbf{c}}_{i_{n}})$. Equation (14) implies that
@@
Thus, we have reduced the event in (11) to the event that
${\operatorname{dist}}({\mathbf{c}}_{i_{1}},H)=O\left(\frac{k^{2}}{(k-l)}\right)\wedge\cdots\wedge{\operatorname{dist}}({\mathbf{c}}_{i_{l}},H)=O\left(\frac{k^{2}}{(k-l)}\right),$
@@
Note that now we have that ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is independent of ${\mathbf{c}}_{j,\{l,\dots,n\}}$ for $i,j\leq l$ and $i\neq j$. However, due to the structure of the Laplacian, ${\mathbf{c}}_{i,\{l,\dots,n\}}$ is not independent of $H_{l,\dots,n}$.
Now, we set $l=k/2$ and we introduce an integer parameter $\tau$, which will be a sufficiently large constant depending on $c$ and $p$, and divide our indices $1,\dots,l$ into $l/2\tau$ sets,
$J_{1}=\{1,\dots,2\tau\},J_{2}=\{2\tau+1,4\tau\},\dots,J_{l/2\tau}=\{l-2\tau+1,\dots,l\}.$
@@
Every coordinate of $\bar{{\mathbf{d}}}_{i,\{l,\dots,n\}}:={\mathbf{d}}_{i,\{l,\dots,n\}}^{*}+{\mathbf{d}}_{i,\{l,\dots,n\}}^{**}$ is an i.i.d. binomial random variable. For a random variable $X\sim B(2\tau,p)$, by Chernoff’s bound, for any $10<\delta<$ 1,
@@
Now, for $i\in[l/2\tau]$, we introduce the following auxiliary randomness inspired again by the “switching” method. We define a new random variable $\hat{{\mathbf{c}}}^{*}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{*})_{j}=({\mathbf{c}}^{*}_{i})_{j}$ for $j\notin D_{i}$, and for $j\in D_{i}$,
$(\hat{{\mathbf{c}}}^{*}_{i})_{j}=\begin{cases}&(\hat{{\mathbf{c}}}_{i}^{*})_{j}\text{ with probability }1/2\\ &({\mathbf{d}}_{i}^{**})_{j}+{\mathbf{f}}_{i}^{*}\text{ with probability }1/2.\end{cases}$
Similarly, we define a new random variable $\hat{{\mathbf{c}}}^{**}_{i}$ where $(\hat{{\mathbf{c}}}_{i}^{**})_{j}=({\mathbf{c}}^{**}_{i})_{j}$ for $j\notin D_{i}$, and for $j\in D_{i}$,
@@
We observe that $\hat{{\mathbf{c}}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$ has the same joint distribution as ${\mathbf{c}}_{i,\{l,\dots,n\}}^{*},H_{l,\dots,n}$. Furthermore, the key benefit of this construction is that, even upon conditioning on the outcome of $H_{l,\dots,n}$, the random vectors ${\mathbf{c}}_{i}$, and ${\mathbf{c}}_{i}^{*}$, $\hat{{\mathbf{c}}}_{i}^{*}$ are sufficiently random with independent entries for $j\in I_{i}$.
@@
On the event $\mathcal{E}$, we have at least $l/2\tau\alpha$ good indices. For any good index $i$, by (15):
${\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{l,\dots,n})\leq O\left(\frac{\tau k^{2}}{(k-l)}\right)\right)$ $\leq{\mathbb{P}}\left({\operatorname{dist}}(\hat{{\mathbf{c}}}^{*}_{i,D_{i}},H_{D_{i}})=O\left(\frac{\tau k^{2}}{(k-l)}\right)\right),$ $\leq C\exp(-\Theta(k))$
where the last inequality follows from Lemma 4.1 and $k$ sufficiently large. Note that our choice of $l$ and the size of $|D_{i}|$ guarantee that the corank of $H_{D_{i}}$ is at least $k/3$. Therefore, by independence of the randomness on the good indices, we have that
@@
Proof of Lemma 4.1.
We show that $f({\mathbf{x}})=\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is convex and 1-Lipschitz. For convexity, we observe that for $0\leq s\leq 1$,
@@
Next, note that for ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{n}$,
@@
${\mathbb{P}}(|f({\mathbf{x}})-M(f({\mathbf{x}})|\geq\lambda T)\leq K\exp(-\kappa\lambda^{2}),$ (18)
@@
$=\left(T^{4}k+T^{2}\|{\mathbf{u}}\|_{2}^{2}k\right),$
@@
Case 1: $(1-2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}\leq\|{\mathbf{u}}\|_{2}\leq(1+2t)T\sqrt{\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}}$. In this case, we have that
@@
By Markov’s inequality, the median of $|Y|$ is $O(T^{2}\sqrt{k})$, which implies that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}^{2}$ is at least $\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}+\|u\|_{2}^{2}-O(T^{2}\sqrt{k})\geq T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})$. We can therefore deduce that the median of $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}$ is at least $\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}-O(T^{2}\sqrt{k})}=\sqrt{T^{2}k+\|{\mathbf{u}}\|_{2}^{2}}-O(T^{2})$. Returning to Talagrand’s inequality, (18), it follows that
@@
Case 2. Now we consider the case where $\|{\mathbf{u}}\|_{2}\leq(1-2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|{\mathbf{u}}\|_{2}\geq(1+2t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$. If $\|P_{H^{\perp}}{\mathbf{x}}-{\mathbf{u}}\|_{2}\leq t\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$, then by the triangle inequality, either $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\leq(1-t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ or $\|P_{H^{\perp}}{\mathbf{x}}\|_{2}\geq(1+t)T\sum_{i=1}^{n}{\mathbf{E}}{\mathbf{x}}_{i}^{2}p_{ii}$ which reduces to Case 1. ∎
5. No-Gaps Delocalization: Proof of Theorem 2.3 for the Approximate Eigenvectors
First, recall the Erdős-Littlewood-Offord result from [55].
@@
  • Assume that at least $N^{\prime}$ elements of the $w_{i}$ satisfy $|w_{i}|\geq r$. Let $X=(x_{1},\dots,x_{N})$ where $x_{i}$ are i.i.d. copies of a random variable $\xi$ of mean zero, variance one, and bounded $(2+\varepsilon)$-moment. Then
@@
  • Furthermore, if we assume that for all $w\in{\mathbb{R}}$, the vector ${\mathbf{w}}-w\mathbf{1}$ has at least $N^{\prime}$ components that have absolute values at least $r$. Then for $X=(x_{1},\dots,x_{N})$ with $x_{i},$ for $1\leq i\leq N$, being i.i.d. copies of $\xi$ we have the following affine analog
@@
$\frac{1}{\log\log n}\leq\lambda_{n}<\lambda_{0},$ (22)
where $\lambda_{0}$ is a sufficiently small positive constant  [^5] This is because the smaller $\lambda_{n}$ is, the stronger our statements become. We postulated this condition on $\lambda_{n}$ primarily due to the union bound in Case 1 below. However, further investigation shows that the events considered in that case for larger $\lambda_{n}$ are, in fact, subevents of those corresponding to smaller $\lambda_{n}$..
It what follows, we will mainly work with $\overline{L}_{n}$, the proof for $\overline{L}^{\prime}_{n}$ (or for $\overline{L}^{\prime}_{n-1}$) is identical.
@@ 5.1. Decomposition of vectors
First, we let $I$ be the set of first $\left\lceil\lambda_{n}n\right\rceil$ indices (in the end we will take union bound over $I$). We will bound the probability of the following event $\mathcal{E}$: there exists a ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) whose first $\lceil\lambda_{n}n\rceil$ coordinates are similar in size (in the $l_{2}$-norm). More precisely, for some $a$
@@
We let $t_{0}>1$ be a parameter whose value is dictated by constraints later in the argument. For instance we can choose
@@
Given ${\mathbf{v}}$, we will partition $[n]$ into two subsets, $I_{m}$ (mixed) and $I_{s}$ (sparse), where $I_{s}$ is the union of the intervals $J_{i_{0}}=I,J_{{i_{1}}},\dots$, where the $i_{k}$ above are defined as follows. At each step $j\geq 0$, find a ${\mathbf{v}}_{J_{i}}$ that has not been selected previously such that there exists $a_{i}\in{\mathbb{Z}}/D$ and a subset $J_{i}^{\prime}\subset J_{i}$ of size at least $|J_{i}|-\alpha_{n}\lambda_{n}n$ which satisfies
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}\leq\delta_{n}t_{0}^{j}.$ (26)
Denote that $i$ as $i_{j}$. This is possible, up to some maximal index $i_{0}$. Hence $I_{s}$ is associated with a maximum index $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ and a partial ordering of $\{1,\dots,\lfloor\lambda_{n}^{-1}\rfloor\}$, where the $J_{i}$ appeared in each step $i$.
@@
Each ${\mathbf{v}}$ will also gives rise to a parameter $0\leq i_{0}\leq 8\lambda_{n}^{-1}$ such that all of the intervals from $I_{s}$ are $\delta_{n}t_{0}^{{i_{0}}}$-close to some $a_{i}\mathbf{1}/\sqrt{n}$. Conversely, if $J_{i}$ does not belong to $I_{s}$ (i.e., if $J_{i}\subset I_{m}$), then by definition, for any $a\in{\mathbb{Z}}/D$ and any $J_{i}^{\prime}\subset J_{i}$ such that $|J_{i}^{\prime}|\geq|J_{i}|-\alpha_{n}\lambda_{n}n$ we have
$\|{\mathbf{v}}_{J_{i}^{\prime}}-(a\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}\|_{2}>\delta_{n}t_{0}^{i_{0}+1}.$
@@
In particular, Condition 21 of Theorem 5.1 is satisfied, and so
$\rho_{L,\frac{\delta_{n}t_{0}^{i_{0}+1}\lambda_{n}^{-1/2}}{\sqrt{n}}}({\mathbf{v}}_{J_{i}})=O\left(\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}}\right).$ (27)
For convenience, set
@@
with an extra factor of $t_{0}^{1/2}$ to be used later.
@@
So, for each fixed $a$, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n}\}$ trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size 1.
As a consequence, the set $\{{\mathbf{v}}_{J_{i}}\in{\mathbb{R}}^{|J_{i}|},\|{\mathbf{v}}_{J_{i}}-a\mathbf{1}/\sqrt{n}\|_{2}\leq t_{0}^{i}\delta_{n},\mbox{ for some }a\in{\mathbb{Z}}/D,|a|\leq\sqrt{n}\}$ trivially accepts an $r_{n}\sqrt{J_{i}}$-net of size $O(D\sqrt{n})$.
We let $I_{m}=[n]\setminus I_{s}$. We write ${\mathbf{v}}_{m}:={\mathbf{v}}_{I_{m}}$ and ${\mathbf{v}}_{s}:={\mathbf{v}}_{I_{s}}$. Note that ${\mathbf{v}}_{s}$ can be $r_{n}\sqrt{|I_{s}|}$-approximated by a vector ${\mathbf{v}}_{s}^{\prime}$, which is a concatenation of vectors of form $(a_{i}\mathbf{1}/\sqrt{n})_{J_{i}^{\prime}}$. We let ${\mathbf{v}}^{\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{m}^{\prime})$, where ${\mathbf{v}}_{m}^{\prime}={\mathbf{v}}_{m}$. Also, based on Subsection 2.4, without loss of generality we can assume the entries of ${\mathbf{v}}_{m}^{\prime}$ are from ${\mathbb{Z}}/D$.
Note that by definition,
@@
where $\sqrt{n}\log n$ comes from the upper bound for $\|L-\lambda_{0}\|_{2}$. By the triangle inequality:
@@
As of this point, roughly speaking, we have extended the special vector ${\mathbf{v}}_{I}$ to a longest possible vector ${\mathbf{v}}_{s}$ whose entries take only a few values, while ${\mathbf{v}}_{m}$ is the left-over that we no longer detect entries with high multiplicities.
Case 1. Let $\mathcal{E}_{1}$ be the event that there exists a non-zero ${\mathbf{v}}\in(\frac{1}{D}{\mathbb{Z}})^{n}$ satisfying (7) and the resulting $I_{m}$ (from ${\mathbf{v}}$) is empty in the sense that for all $i$ we can approximate ${\mathbf{v}}_{J_{i}}$.
@@
We will consider a $\delta_{n}t_{0}^{8\lambda_{n}^{-1}}$-approximation of the vectors. There are at most $D^{\alpha_{n}n}\times D^{8\lambda_{n}^{-1}}$ choices of ${\mathbf{v}}^{\prime}$ that can arise with empty $I_{m}$. The first factor stands for the amount of choices for the outliers; the second factor stands for the possible choices of the representative $a_{i}$ of each interval, assuming that $1/D$ is small compared to $\delta_{n}t_{0}^{\lambda_{n}^{-1}}$, which is necessary for the approximation to make sense.
@@
Then either the smallest or largest values of $a_{i}$ is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the rest. Assume without loss of generality that $a$, the multiplicity of the first segment, is $\delta_{n}^{\prime}$-far from $\lambda_{n}^{-1}/2$ of the multiplicities $j_{0}$.
Consider the event